Roll two fair six-sided dice, and let X, Y denote the first and the second numbers.


If Z=max {X, Y}, find

- E(Z)

- V(Z)

If Z=|X-Y|, find

- E(Z)

- V(Z)

Roll two fair six-sided dice, and let X, Y denote the first and the second numbers. If Z max X, Y), find . E(Z) . V(Z) . E(Z) V(Z)
Answer #1

The possible values of (X,Y) are

{(1,1),(1,2),(1,3),(1,4),(1,5),(1,6),

(2,1),(2,2),(2,3),(2,4),(2,5),(2,6),

(3,1),(3,2),(3,3),(3,4),(3,5),(3,6),

(4,1),(4,2),(4,3),(4,4),(4,5),(4,6),

(5,1),(5,2),(5,3),(5,4),(5,5),(5,6),

(6,1),(6,2),(6,3),(6,4),(6,5),(6,6)}

Thus, Distribution of Z = max{X,Y} is

Z 1 2 3 4 5 6
P(Z) 1/36 3/36 5/36 7/36 9/36

11/36

Thus, E(Z) = sum z.P(z)

= 161/36

= 4.472

V(Z) = E(Z^{2})-E(Z)^{2}

= 2. P(z) 4.47 2

= 791/36 - 20

= 1.972

Now, Distribution of Z = I X - Y I is

Z 0 1 2 3 4 5
P(Z) 6/36 10/36 8/36 6/36 4/36 2/36

hus, E(Z) = sum z.P(z)

= 70/36

= 1.9444

V(Z) = E(Z^{2})-E(Z)^{2}

= 2P(z) - 1.944

= 2.0525